# Towards General Industrial Intelligence: A Survey on Industrial IoT-Enhanced Continual Large Models

### v1

@@ IV-A Classical Continual Learning
Continual LearningReplay-basedRegularization-basedParameter IsolationParameter RegularizationFunctional RegularizationExperience ReplayGenerative ReplayStatic ArchitectureDynamic Architecture EWC [129], Online EWC [130], SI [131], MAS [132], TFCL ([133, 134, 135]), SPLR [136], KCL [137], ACML [138], RTRA [139], CoSR [140] LwF[141], ANCL [142], TAM-CL [143], PRD [144], TARGET [145] iCaRL [146], TinyER [147], DER [148], EPR [149], AdaER [150], GEM [151], A-GEM [152], OCD-Net [153], CRL [154], PODNet [155], DER/DER++ [148], D3former [156], AAER [157], EPR [158], CPL [159], RR [160], InfluenceCL [161], BFP [162], LP-DiF [163] DGR [164], DGM [165], FearNet [166], SRM [167], EEC [168], MAE [169], DDGR [170], CPDM [171], DiffusePast [172], FSIL-GAN [173] DSG [174] PathNet [175], PackNet [176], HAT [177], RPS-Net [178], UCL [179], CLNP [180], CDG [181], DSN [182] PNNs [183], Expert Gate [184], CCLL [185], PI-GNN [186], DEGC [187], SEDEM [188], SANE [189], Dynamos [190], Expo [191]
Figure 23: Taxonomy of CL.
IV-A1 Regularization-based Methods
@@
Regularization-based methods commonly employ regularization terms to prevent catastrophic forgetting during training. According to the regularization target, these methods can be further categorized into two sub-directions: parameter regularization and functional regularization.
@@
Parameter Regularization Methods. In parameter regularization methods, regularization terms constrain the variations of parameters to balance the plasticity-stability trade-off. The regularization term is calculated to discourage significant changes in these critical parameters while training new tasks, as shown in Fig. 24. For task $t$ with the parameter $\theta$, the loss function ~$\tilde{\mathcal{L}}_{t}(\theta)$ is [131]
@@
where $\mathcal{L}_{t}(\theta)$ is the task-specific loss for task $t$, $i$ labels each parameter, $\Omega^{t}_{i}$ is the importance matrix of the $i$-th parameter when learning task $t$, $\theta_{t-1,i}^{*}$ is the $i$-th parameter after learning task $t-1$, and $c$ is a hyperparameter that determines the importance of the previous task compared to the new one. As the first work proposing the regularization-based method, Elastic Weight Consolidation (EWC) [129] estimates the importance of each parameter based on its significance for the previous tasks. Consider the training from a Bayesian framework [192], the conditional probability $p\left({\theta}\mid\mathcal{D}_{1:t}\right)$ after learning the task $t$ can be obtained by [193]
$\centering\begin{split}p\left({\theta}\mid\mathcal{D}_{1:t}\right)&\propto p({\theta})\prod_{i=1}^{t}p\left(\mathcal{D}_{i}\mid{\theta}\right)\\ &\propto p\left({\theta}\mid\mathcal{D}_{1:t-1}\right)p\left(\mathcal{D}_{t}\mid{\theta}\right),\end{split}\@add@centering$ (10)
where the distribution $p\left({\theta}\mid\mathcal{D}_{1:t-1}\right)$ incorporates all information from previous tasks [129]. Typically, the posterior distribution $p\left({\theta}\mid\mathcal{D}_{1:t-1}\right)$ is intractable, necessitating approximation (e.g., Laplace approximation, Variational inference, and Moment matching). In Laplace approximation [192], the posterior distribution is treated as a multivariate Gaussian, i.e., $p(\theta\mid\mathcal{D}_{1:t-1})\approx q_{t-1}(\theta)\approx\mathcal{N}\left(\theta;\theta_{t-1}^{*},\Lambda_{t-1}^{-1}\right)$, where $\Lambda_{t-1}$ is the precision matrix. To calculate the precision matrix, EWC applies Fisher information matrix (FIM) with a diagonal approximation to approximate the Hessian matrix. Then, the loss function when learning task $t$ in EWC becomes
~$\centering\tilde{\mathcal{L}}_{t}(\theta)=\mathcal{L}_{t}(\theta)+c\sum_{i}F_{i}\left(\theta_{i}-\theta_{t-1,i}^{*}\right)^{2},\@add@centering$ (11)
@@
Besides Laplace approximation, Variational inference (VI) [193, 194] is also commonly employed for approximating the posterior distribution. VI utilizes optimization techniques to estimate the intractable probability distributions. It aims to determine the distribution that best approximates the actual one, typically by using the KL divergence as the optimization criterion [195]:
$\centering q_{t}({\theta})=\arg\min_{q\in\mathcal{Q}}\operatorname{KL}\left(q({\theta})\|\frac{1}{Z_{t}}q_{t-1}({\theta})p\left(\mathcal{D}_{t}\mid{\theta}\right)\right),\@add@centering$ (12)
where $\mathcal{Q}$ is the set of allowed approximate posteriors, $q_{0}(\theta)=p(\theta)$, and $Z_{t}$ is a normalizing constant. Variational Continual Learning (VCL) [195] combines online variational inference (VI) with Monte Carlo VI [196] to construct a general framework for both deep discriminative models and deep generative models. The VI formulation naturally includes a regularization term. In VCL, training in Eq. 12 is equivalent to maximizing the variational lower bound to the negative online variational free energy or the online marginal likelihood, which is
$\centering\mathcal{L}_{t}\left(q_{t}({\theta})\right)=\mathbb{E}_{{\theta}\sim q_{t}({\theta})}(\ell_{t}(\theta))-\mathrm{KL}\left(q_{t}({\theta})\|q_{t-1}({\theta})\right),\@add@centering$ (13)
where the KL-divergence can be computed in closed form.
Figure 24: Illustration of the effect of parameter regularization across two tasks: $\theta_{1}$ represents the parameter after learning task 1. $\theta_{2}$, $\dot{\theta}_{2}$, and $\ddot{\theta}_{2}$ represent the parameters after learning task 2 under three conditions: parameter regularization, $L_{2}$ regularization, and no regularization, respectively. While $\theta_{2}$ learns task 2 while still retaining knowledge of task 1, $\dot{\theta}_{2}$ learns only task 1 under the overly stringent constraint, and $\ddot{\theta}_{2}$ learns only task 2, forgetting task 1 without any penalty [129].
Considerable efforts have been committed to enhancing parameter regularization approaches. SI [131] performs the importance measurement by considering the parameter’s contribution to the loss decay in an online manner and along the entire learning trajectory. MAS [132] accumulates the importance measure for each parameter via the sensitivity of the predicted output function to changes based on unlabeled data in an online manner. TFCL [133] combine MAS with online CL to extend task-based sequential CL into task-free settings, where data distributions change gradually and the concept of distinct tasks is absent. SPLR [136] learns to task invariant representation and task-specific features across the entire parameter trajectory via SI.
For those methods that employ VI strategies, building upon VCL, CLAW [197] adapts the local and global components of the model architecture with an attention mechanism. For each neuron, one binary parameter controls whether adaptation occurs, along with two parameters regulating the magnitude of adaptation. KCL [137] utilizes an episodic memory unit to store a subset of samples for each task, enabling the learning of task-specific classifiers through kernel ridge regression. ACML [138] establishes a one-to-many mapping between a task and meta-knowledge components to facilitate meta-knowledge sharing among heterogeneous tasks through an Indian Buffet Process prior [198].
Functional Regularization Methods. Most functional regularization methods are based on knowledge distillation. Their key strategy to prevent catastrophic forgetting is recording the network outputs of a new task before its training, then utilizing these recorded outputs to distil the previously learned knowledge [199]. One classical functional regularization method is LwF [141]. The learning procedure for each task in LwF is divided into two stages: initialization and training. During the initialization stage, the output of the previous networks for the new task is computed and recorded, and the task-specific parameters for a new task are randomly initialized. During the training stage, two losses are utilized to update parameters: one for new specific tasks ($\mathcal{L}_{\text{n}}$) and another for distillation ($\mathcal{L}_{\text{o}}$). For new tasks, the loss $\mathcal{L}_{\text{n}}$ is calculated based on the ground truth $\mathbf{y}_{n}$ and the predictions $\hat{\mathbf{y}}_{n}$. For each original task, the knowledge distillation loss $\mathcal{L}_{\text{o}}$ is applied to ensure that the output probabilities closely align with the recorded outputs from the original network. As for multi-class classification with $l$ labels, these two loss can be estimated by
$\centering\mathcal{L}_{\text{n}}\left(\mathbf{y}_{n},\hat{\mathbf{y}}_{n}\right)=-\mathbf{y}_{n}\log\hat{\mathbf{y}}_{n},\@add@centering$ (14)
$\centering\mathcal{L}_{\text{o}}\left(\mathbf{y}_{o},\hat{\mathbf{y}}_{o}\right)=-\sum_{i=1}^{l}y_{o}^{\prime(i)}\log\hat{y}_{o}^{\prime(i)},\@add@centering$ (15)
where $y_{o}^{\prime(i)}$ and $\hat{y}_{o}^{\prime(i)}$ are modified by softmax temperature ($T=2$ in LwF) [200] from the recorded probabilities $y_{o}^{(i)}$ and current probabilities $\hat{y}_{o}^{(i)}$. Thus, the optimization objective for the shared parameters $\theta_{s}$, task-specific parameters for previous tasks $\theta_{o}$, and the new task-specific parameters $\theta_{n}$ is as follows:
[…]
where $\lambda$ is a hyperparameter that determines the weight of the previous task performance and $\mathcal{R}$ is a regularizer.
A growing number of functional regularization variants have been developed. EBLL [201] constructs a task-specific under complete autoencoder for each task to capture the critical features from previous tasks. It then ensures that the reconstruction of these features from changing is prevented when learning a new task. Based on LwF, LwM [202] proposes an attention-based method to transfer knowledge and restrict the divergence between student and teacher in CIL problems. DMC [203] introduces a double distillation loss to mitigate the bias caused by information asymmetry between previous and new classes. AYT [204] introduces Teacher Adaptation, allowing teacher and student models to update simultaneously when learning new tasks in exemplar-free CIL settings.
IV-A2 Replay-based Methods
@@
Replay-based methods prevent catastrophic forgetting by replaying a part of data from previous tasks and using them when learning a new task. Depending on the way to replay the data, replay-based methods can be further categorized into two classes: experience replay/rehearsal and generative replay/pseudo-rehearsal.
Experience Replay Methods. In experience replay methods, the model trains using a part of previous data samples stored in a memory buffer B and the new task data to prevent catastrophic forgetting (see Fig. 25). Based on LwF, iCaRL [146] proposes this idea to tackle CIL problems. It sets the exemplar sets $\mathcal{P}$ that updates whenever learning new classes. Here, a prioritized list $P$ represents the exemplar set from a class, and $p$ represents a specific exemplar within $P$. $P$ is defined as $P=\left(p_{1},p_{2},...,p_{m}\right)$. The examples at the beginning of the set are more important than those that follow. $m$ denotes the capacity of an exemplar set, i.e., the number of exemplars it can store. If $K$ denotes the total number of exemplars that can be stored and $t$ is the number of observed classes, then $m$ can be calculated as $m=K/t$. For $i=1,2,...,m$ in the exemplar set selection, the exemplar $p_{i}$ is selected by
@@
$\centering p_{i}\leftarrow\underset{x\in X}{\operatorname{argmin}}\left\|\mu-\frac{1}{i}\left[\varphi(x)+\sum_{j=1}^{i-1}\varphi\left(p_{j}\right)\right]\right\|,\@add@centering$ (17)
where $X=\left\{x_{1},x_{2},...,x_{n}\right\}$ denotes the data set of a class, $\varphi(\cdot)$ is the feature function, and $\mu$ is the class mean that calculated by $\mu=\frac{1}{n}\sum_{x\in X}\varphi(x)$. Upon starting to learn class $t$, iCaRL possesses the exemplar set $\mathcal{P}=\left(P_{1},P_{2},...,P_{t-1}\right)$, containing the exemplars selected from class $1$ to class $t-1$.
In addition to using herding selection as iCaRL does [205, 155], various methods can be used, e.g., reservoir sampling [147, 148, 153, 206], gradient-based sampling [207], ring buffer [151], k-Means [147], etc. Moreover, various methods for enhancing sampling strategies are proposed [208, 209, 149, 150]. EPR [149] selects the previous data through saliency maps that provide visual explanations for the model predictions. AdaER [150] selects examples based on contextual cues that consider the interference caused by the data-conflicting and the task performance.
Instead of improving sampling strategies, some methods prioritize learning more knowledge from the memory buffer. GEM [151] treats the loss for each previous task as inequality constraints to prevent their increase, and violations are addressed using gradient projection. Some advancements apply experience replay combined with other machine learning methods, such as meta-learning [210] and contrastive learning [211, 212, 153, 154, 206, 213]. KD is also widely employed in experience replay [155, 148, 156, 157, 158, 159, 160, 161, 163]. PODNet [155] utilizes spatial-based distillation loss all over the network to constrain representation evolution. AAER [157] utilizes anchors to tackle drift prevention and compensation. BFP [162] introduces a new distillation loss based on feature projection to preserve the linear separability of previous classes while maintaining the plasticity.
Figure 25: Framework of two replay-based methods for learning task $t$. In the experience replay method, the new continual model $M_{t}$ is trained on both the current input $(x_{t},y_{t})$ from task $t$ and the replayed data $(x_{t}^{\prime},y_{t}^{\prime})$. In the generative replay method, the generated input $x_{t}^{\prime}$ is input into the previous continual model $M_{t-1}$. The output $y_{t}^{\prime}$ from $M_{t-1}$, paired with $x_{t}^{\prime}$, is utilized to train the new continual model $M_{t}$ with $(x_{t},y_{t})$. The new generator $G_{t}$ is trained on $x_{t}$ and $x_{t}^{\prime}$.
Generative replay Methods. Generative replay methods train an additional generative model to generate pseudo-data for previous tasks, while experience replay methods require substantial memory resources to store and replay previous data, rendering them potentially impractical in specific real-world scenarios [164, 174]. DGR [164] utilizes GANs as a generator to replay previous data and represents the model $H$ as $\left\langle G,S\right\rangle$, where $G$ is the generator and $S$ is a task solving model. The scholar’s training involves training the generator $G$ and training the solver $S$. The generator is trained on the current input $\bm{x}$ and the replayed input $\bm{x}^{\prime}$. The solver is trained on the current data pair $\left(\bm{x},y\right)$ and the replayed data pairs $\left(\bm{x}^{\prime},y^{\prime}\right)$, where $y^{\prime}$ is obtained by inputting $\bm{x}^{\prime}$ into the old solver. When learning the task $t$, the scholar $H_{t}$ learns both the task $t$ and the knowledge of $H_{t-1}$. The train loss is computed by
$\mathcal{L}_{t}(\theta)$ $=r\mathbb{E}_{(\bm{x},\bm{y})\sim D_{t}}[L(S_{t}(\bm{x}),\bm{y})]$ (18)
$\quad+(1-r)\mathbb{E}_{\bm{x}^{\prime}\sim G_{t-1}}[L(S_{t}(\bm{x}^{\prime}),S_{t-1}(\bm{x}^{\prime}))],$
where $r$ is the mixing ratio of real and replayed samples.
Many methods use GANs as generators [214, 215]. DGM [165] introduces conditional generative networks that learn sparse attention masking for connection weights or layer activations during adversarial training. GFR [216] integrates GANs with a feature extractor for feature replay. GNR [217] utilizes the generated data as negative examples to improve the learning of current class features. GarDA [218] and MFA [219] investigate the utilization of GANs in unsupervised domain adaptation [220].
Autoencoders and VAEs [221] are also utilized for data generation. Earlier approaches include but are not limited to FearNet [166], SRM [167], and EEC [168]. MAE [169] employs masked autoencoders that randomly mask patches of input images for enhancing replay efficiency and stability. More recently, diffusion probabilistic models [222] have been employed as generators [170, 171]. KD is applied to generative replay as well [218, 219, 223, 224].
IV-A3 Parameter Isolation Methods

### v2

@@ VI-A Theoretical Insights
As shown in Fig. 24, CL is a strategy for training machine learning models to adapt to a sequence of changing tasks, each potentially introducing new data distributions [200, 201]. We define a series of tasks $\{\mathcal{D}_{t}\}_{t=1}^{T}$, where each task $\mathcal{D}_{t}$ consists of a set of data samples and corresponding labels, i.e., $\mathcal{D}_{t}=\{(\mathbf{x}_{i}^{t},y_{i}^{t})\}_{i=1}^{|\mathcal{D}_{t}|}$, where $\mathbf{x}_{i}^{t}$ is the input data and $y_{i}^{t}$ is the corresponding label. The goal in each task stage $t$ is for the model to learn the new task’s data while addressing catastrophic forgetting on previous tasks.
Class-Incremental Learning (CIL) is an important branch of CL. In CIL, each task $\mathcal{D}_{t}$ introduces an independent set of categories, i.e., $y_{i}^{t}\in\mathcal{Y}_{t}$ and $\mathcal{Y}_{t}\cap\mathcal{Y}_{j}=\emptyset$ for all $t\neq j$. According to the goals of CL, the objective for CIL is to fit a model $f(\cdot)$, and minimize the cumulative expected risk across the entire sequence of tasks:
$f^{*}=\arg\min_{f\in\mathcal{H}}\sum_{t=1}^{T}\mathbb{E}_{(\mathbf{x}_{i}^{t},y_{i}^{t})\sim\mathcal{D}_{t}}\left[\mathbb{I}[f(\mathbf{x}_{i}^{t})\neq y_{i}^{t}]\right],$ (9)
where $\mathcal{H}$ is the hypothesis space, $\mathbb{I}(\cdot)$ is the indicator function which outputs 1 if the expression holds and 0 otherwise.
Two widely used metrics in continual learning are Average Accuracy (ACC) and Forgetting Rate:
  • ACC: This metric evaluates the model’s overall performance across all tasks after training is completed. It is calculated as the average of the accuracies achieved on all tasks, providing an indication of how well the model retains knowledge of previous tasks while learning new ones.
  • Forgetting Rate: This metric quantifies the degree of performance degradation on earlier tasks as new tasks are learned. It is computed as the difference between the maximum accuracy achieved on a task and the final accuracy on the same task. A lower forgetting rate indicates better retention of past knowledge.
VI-B IIoT for Continual Learning
The IIoT serves not only as an application domain for continual learning but also as a critical enabling platform. By providing dynamic data streams, edge computing capabilities, and efficient communication networks, IIoT establishes the essential infrastructure for model training, deployment, and real-time updates in continual learning. For example, in the integration of FL and edge computing (MEC), the distributed architecture of IIoT enables cross-device collaboration, supporting knowledge sharing and incremental learning across heterogeneous devices, thereby enhancing model adaptability and efficiency.
Specifically, research in this area focuses on two major directions: the integration of FL with edge computing and the application of continual learning in specific industrial scenarios.
Integration of FL with Edge Computing. IIoT’s distributed environment provides a natural foundation for federated learning across devices. Frameworks such as Cross-FCL [202] utilize parameter decomposition, global aggregation, and local optimization strategies to retain knowledge and task memory across multiple edge devices, improving model efficiency and accuracy. Additionally, personalized federated learning frameworks [203, 204] are deployed on IIoT devices to enable incremental task learning while reducing communication costs. These methods fully leverage the heterogeneity and distributed nature of IIoT to better adapt models to dynamic environments.
Applications in Specific Industrial Scenarios. IIoT’s dynamic data support enables a wide range of applications for continual learning in specific industrial scenarios. For instance, in digital twin synchronization, IIoT provides real-time sensor data streams to maintain synchronization between physical objects and their digital representations [205], using incremental training to handle data drift and real-time updates. In electromagnetic signal classification, methods like SMTC-CL [206] address task increments and catastrophic forgetting through selective multi-task coordination. Similarly, in QoE assessment, deep learning-based models leverage continuously incoming data from IIoT to enable efficient evaluation while reducing training complexity [207].
VI-C Traditional Continuous Learning
Continual LearningReplay-basedRegularization-basedParameter IsolationParameter RegularizationFunctional RegularizationExperience ReplayGenerative ReplayStatic ArchitectureDynamic Architecture EWC [208], Online EWC [209], SI [210], MAS [211] LwF[212], DMC [213], AYT [214] iCaRL [215], GEM [216], A-GEM [217], OML [218] DGR [219], ESGR [220], DDGR [221], DSG [222], PathNet [223], PackNet [224], HAT [225] PNNs [226], DEN [227]
Figure 25: Taxonomy of Traditional CL.
As shown in Fig. 25, existing continual learning surveys generally classify methods into the following categories:
  • Regularization-based methods: These mitigate catastrophic forgetting of previous tasks by adding regularization terms to the loss function.
  • Replay-based methods: These alleviate catastrophic forgetting within a limited memory buffer by using data from previous tasks.
  • Architecture-based methods: These allocate independent parameters for new task data, for example, through parameter allocation, model partitioning, or modular network models.
VI-C1 Regularization-Based Methods
@@
Regularization-based methods aim to mitigate catastrophic forgetting by incorporating regularization terms into the training process. These methods can be broadly categorized into two subtypes: parameter regularization and functional regularization.
Parameter Regularization Methods.
Figure 26: Illustration of the effect of parameter regularization across two tasks.
@@
Parameter regularization constrains changes to critical parameters, balancing the plasticity-stability trade-off. The regularization term penalizes significant deviations in important parameters during new task training, as illustrated in Fig. 26. For a task $t$, given the model parameters $\theta$, the regularized loss ~$\tilde{\mathcal{L}}_{t}(\theta)$ is expressed as [210]:
@@
where $\mathcal{L}_{t}(\theta)$ represents the task-specific loss for task $t$, $\Omega_{i}^{t}$ is the importance matrix of the $i$-th parameter for task $t$, $\theta_{t-1,i}^{*}$ is the $i$-th parameter value after task $t-1$, and $c$ is a hyperparameter controlling the balance between stability and plasticity.
Elastic Weight Consolidation (EWC) [208] was the first method to use parameter regularization, estimating parameter importance using the Fisher Information Matrix (FIM). From a Bayesian perspective, the conditional probability $p(\theta\mid\mathcal{D}_{1:t})$ for parameters after task $t$ can be expressed as:
$p(\theta\mid\mathcal{D}_{1:t})\propto p(\theta)\prod_{i=1}^{t}p(\mathcal{D}_{i}\mid\theta)\propto p(\theta\mid\mathcal{D}_{1:t-1})p(\mathcal{D}_{t}\mid\theta),$ (11)
where $p(\theta\mid\mathcal{D}_{1:t-1})$ encapsulates information from previous tasks [208]. However, this posterior distribution is often intractable and requires approximations such as Laplace approximation, variational inference, or moment matching. Using Laplace approximation, the posterior is approximated as a multivariate Gaussian:
$p(\theta\mid\mathcal{D}_{1:t-1})\approx\mathcal{N}(\theta;\theta_{t-1}^{*},\Lambda_{t-1}^{-1}),$ (12)
where $\Lambda_{t-1}$ is the precision matrix. EWC approximates $\Lambda_{t-1}$ using a diagonal FIM, resulting in the following regularized loss:
~$\tilde{\mathcal{L}}_{t}(\theta)=\mathcal{L}_{t}(\theta)+c\sum_{i}F_{i}\left(\theta_{i}-\theta_{t-1,i}^{*}\right)^{2},$ (13)
@@
Subsequent efforts have refined parameter regularization. SI [210] measures parameter importance based on its contribution to the loss decay along the training trajectory, while MAS [211] calculates importance via the sensitivity of output predictions to parameter changes using unlabeled data.
Functional Regularization Methods. Functional regularization primarily leverages knowledge distillation to retain knowledge from prior tasks. The central idea is to preserve the network’s output predictions for previous tasks while training on new tasks [228]. Learning without Forgetting (LwF) [212] serves as a foundational method in this category. It achieves this by employing two loss components: one to guide task-specific learning for the new task and another to align the current predictions with those generated by the original network on prior tasks. This ensures that the model retains its ability to perform previous tasks without catastrophic forgetting.
Instead of relying on detailed equations, LwF utilizes a distillation-based approach, where the network’s softened probabilities (calculated using a temperature scaling factor) are used to balance the performance on both old and new tasks. The relative importance of these components is controlled through weighting factors, which ensure that past knowledge is preserved without hindering the learning of new tasks.
Recent advancements in functional regularization have extended its capabilities. For instance, DMC [213] incorporates double distillation techniques to mitigate biases between new and old classes, thereby improving class-incremental learning performance. Similarly, AYT [214] enables synchronous updates of teacher and student models, making it particularly effective in exemplar-free class-incremental learning scenarios. These methods reflect the continuous evolution of functional regularization strategies to better address the challenges of continual learning.
VI-C2 Replay-Based Methods
@@
Replay-based methods mitigate catastrophic forgetting by replaying data from previous tasks during the learning of new tasks. Depending on how data is replayed, these methods are categorized into two main types: experience replay and generative replay.
Experience Replay Methods. Experience replay methods store a subset of data samples from previous tasks in a memory buffer B and combine them with data from the current task during training. This approach effectively preserves knowledge from earlier tasks and prevents catastrophic forgetting.
Based on Learning without Forgetting (LwF), iCaRL [215] introduces an exemplar-based memory mechanism for class-incremental learning (CIL). It maintains exemplar sets $\mathcal{P}$, which are updated whenever new classes are learned. Each exemplar set $P$ for a specific class is defined as $P=\left(p_{1},p_{2},...,p_{m}\right)$, where $m$ denotes the capacity of the set. The total memory capacity $K$ is distributed across all observed classes $t$, such that $m=K/t$.
The selection of exemplars is performed iteratively. For the $i$-th exemplar in the set, $p_{i}$ is chosen as:
@@
$p_{i}\leftarrow\underset{x\in X}{\operatorname{argmin}}\left\|\mu-\frac{1}{i}\left[\varphi(x)+\sum_{j=1}^{i-1}\varphi\left(p_{j}\right)\right]\right\|,$ (14)
where $X=\{x_{1},x_{2},...,x_{n}\}$ represents the data samples for a class, $\varphi(\cdot)$ is the feature function, and $\mu$ is the class mean, computed as:
$\mu=\frac{1}{n}\sum_{x\in X}\varphi(x).$ (15)
At the start of training for class $t$, the memory buffer $\mathcal{P}=\{P_{1},P_{2},...,P_{t-1}\}$ stores exemplars from previously observed classes.
Beyond sampling strategies, other methods focus on effectively leveraging stored data. Gradient Episodic Memory (GEM) [216] imposes inequality constraints on the losses of previous tasks to avoid forgetting, addressing violations through gradient projection. OML [218], a meta-learning approach, trains a deep Representation Learning Network (RLN) to learn sparse and robust features, enhancing its ability to retain knowledge across tasks.
Generative Replay Methods. Generative replay methods employ a generative model to synthesize pseudo-data that represents previous tasks, thereby reducing the dependency on large memory buffers typically required in experience replay [219, 222]. As illustrated in Fig. 14, there are three common types of generative models used in generative replay: Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and Diffusion Models.
GANs have been extensively applied in generative replay due to their ability to produce high-quality and realistic data [220, 219]. Recently, diffusion models [221, 222] have also been utilized for data generation, leveraging their robustness and capability to capture complex data distributions. These methods effectively mitigate memory constraints and ensure knowledge retention, making them particularly suitable for real-world continual learning applications.
VI-C3 Parameter Isolation Methods